alice is twice as old as brian. cristal is three years younger then brian .the sum of their ages is 45. how old is cristal?
Answer by edjones(8007)

Let a=alice's age, b=brian's and c=crystal's
a=2b
c=b-3
.
b+2b+b-3=45
4b-3=45
4b=48
b=12
a=24
c=9
.
Ed